OpenStudy (jojo12g):

Which property is shown? (–9 + 17) + 23 = –9 + (17 + 23) A. associative property B. opposite of a sum property C. commutative property D. distributive property

OpenStudy (vjt):

Do you know what the different properties do?

OpenStudy (jaelenecastro1998):

I believe its commutative property

OpenStudy (vjt):

The associative property states that you can add or multiply regardless of how the numbers are grouped. By 'grouped' we mean 'how you use parenthesis'. In other words, if you are adding or multiplying it does not matter where you put the parenthesis. Add some parenthesis any where you like. Do you understand this?

OpenStudy (jaelenecastro1998):

Never mind its associative

umerlodhi (umerlodhi):

sorry i was busy im thinkin A

OpenStudy (jaelenecastro1998):

your right @VJT

OpenStudy (vjt):

sorry accidentally deleted that. Pretty much said that its not the commutative property because with that you flip the equation so that you will end up with the same answer on each side so this makes it associative.

OpenStudy (skullpatrol):

$$\Huge ( –9 + 17) + 23 = –9 + (17 + 23)$$ $$\Huge (a + b) + c = a + ( b + c) $$

OpenStudy (skullpatrol):

The association of the numbers does not affect the result. In this example on the LHS ( 8 ) + 23 = 31. While on the RHS –9 + ( 40 ) = 31 The result of 31 is not affected by the association of the numbers i.e. how they are grouped